# Flow Matching Meets PDEs: A Unified Framework for Physics-Constrained Generation

### v1

@@ Abstract
Generative machine learning methods, such as diffusion models and flow matching, have shown great potential in modeling complex system behaviors and building efficient surrogate models. However, these methods typically learn the underlying physics implicitly from data. We propose Physics-Based Flow Matching (PBFM), a novel generative framework that explicitly embeds physical constraints, both PDE residuals and algebraic relations, into the flow matching objective. We also introduce temporal unrolling at training time that improves the accuracy of the final, noise-free sample prediction. Our method jointly minimizes the flow matching loss and the physics-based residual loss without requiring hyperparameter tuning of their relative weights. Additionally, we analyze the role of the minimum noise level, $\sigma_{\min}$, in the context of physical constraints and evaluate a stochastic sampling strategy that helps to reduce physical residuals. Through extensive benchmarks on three representative PDE problems, we show that our approach yields up to an 8×8$\times$8 × more accurate physical residuals compared to FM, while clearly outperforming existing algorithms in terms of distributional accuracy. PBFM thus provides a principled and efficient framework for surrogate modeling, uncertainty quantification, and accelerated simulation in physics and engineering applications.
@@ 2.3 Improving physical and distributional accuracy
An important requirement for generative models in the context of physics applications is the necessity to produce high-accuracy samples that are on-par with traditional methods. We identify the reconstruction of the final, noise-free, sample during training as a crucial aspect for accurately evaluating the physical residual. The integration of the intermediate sample at time $t$ forward to the final time is performed with ODE integration (Lipman et al., 2024). Employing the predicted velocity field $u_{t}^{\theta}$ by integrating over a single time step of size $1-t$. While trajectories are straight in theory, relying on just one integration step limits prediction accuracy, particularly when $t$ is close to zero. To improve the final prediction, we incorporate unrolling (Monga et al., 2021), by performing $n$ ODE integration steps of size $(1-t)/n$. The integration over multiple time steps has an averaging effect, which naturally promotes trajectories that are closer to the theoretically straight paths. Unrolling is applied via a curriculum, increasing the number of unrolling steps over the course of training. We additionally down-weigh less accurate predictions near $t=0$ by applying a scaling factor of $t^{p},\ p_{\text{opt}}=1$ in the residual loss computation. A detailed analysis of different power laws is provided in Appendix D. The more accurate prediction of $x_{1}$ substantially improves the evaluation of the physics residual in eq. 2, which in turn yields a better learning direction.
@@
$dt\leftarrow(1-t)/n$triangleright Compute $dt$ to reach the final state at $t=1$
$\widetilde{t}\leftarrow t$triangleright Keep the starting time $t$ to weight the residual loss
$u_{t}^{\theta}\leftarrow\text{model}(x_{t},\ t)$triangleright Flow matching velocity $u_{t}^{\theta}$ is not part of unrolling
@@
     $\widetilde{u}_{t}^{\theta}\leftarrow\text{model}(\widetilde{x}_{1},\ \widetilde{t})$
     $\widetilde{x}_{1}\leftarrow\widetilde{x}_{1}+dt\cdot\widetilde{u}_{t}^{\theta}$triangleright Current fields $\widetilde{x}_{1}$ are updated until $t=1$
@@
$\mathcal{R}\leftarrow$ compute residual$(\widetilde{x}_{1})$
$\mathcal{L}_{\mathcal{R}}\leftarrow\|t^{p}\cdot\mathcal{R}\|_{2}$triangleright Residuals are weighted with $t^{p},\ p_{\text{opt}}=1$
@@
$\nabla_{\theta}$leftarrow compute $\mathbf{g}_{\text{update}}$ via Eq. 3 $\triangleright$ Conflict-free update
@@ 3 Experimental Setup
The second benchmark is the two-dimensional Kolmogorov flow over a $128\times 128$ spatial domain with periodic boundary conditions. The dataset is generated using a spectral solver and consists of velocity field snapshots for various Reynolds numbers in the range [100; 500]100500[100;\ 500][ 100 ; 500 ]. The Reynolds number conditions the data generation, influencing the turbulence scales and flow complexity. The training set includes data for 32 Reynolds numbers with 1024 temporal snapshots each. The test set additionally includes 16 unseen Reynolds numbers. The data distribution reflects the temporal variation within each flow regime. The prediction target consists of the two velocity components, which are expected to satisfy the conservation of mass via:
@@ 4.1 Darcy flow
We begin by evaluating overall performance, focusing on residual mean absolute error (MAE), along with pressure and permeability distributions. The results are summarized in Figure 3. The first panel illustrates how residual error evolves with the number of unrolling steps used during training. The bar plot shows that MAE decreases significantly from approximately 1.7 for FM to about 0.6 for PBFM. Notably, the improvement achieved by introducing the additional loss term is comparable in magnitude to the improvement gained via unrolling. When comparing across methods, we observe that the DM baseline performs similarly to FM. CoCoGen produces an $\approx\!10\%$ lower residual error than FM but ca. $20\%$ higher than PBFM-1, at the same time incurring a higher compute cost for inference. While PIDM-ME yields the lowest residual error, this performance comes at a steep cost: As shown in the last two panels, the residual scores for PIDM-ME are associated with a severely degraded pressure distribution, characterized by values constrained within the narrow range ±$\pm 0.2$. This artificially reduces the residual due to the lower magnitude of $P$, rather than improved fidelity. In comparison, all variants of our framework more faithfully reproduce the true pressure distribution. Regarding permeability, the DM baseline, as well as CoCoGen, exhibit a loss of distributional fidelity, particularly at the extrema. While our framework also shows a slight degradation for the least probable values, it successfully recovers the maximum $K$ values around 30, which are approximately six orders of magnitude less frequent than the central mode of the distribution. Figure 4 visualizes the trend of decreasing residual error as the FM model is progressively improved.
@@
To complete the analysis on the Darcy dataset, we evaluated the PBFM-1 framework for sampling time $t$ during training with the logit-normal distribution, comparing it to the standard uniform distribution. We also analyze the effect of different samplers: deterministic and stochastic. Figure 5 reports the residual error along with the Wasserstein distance (Panaretos and Zemel, 2019; SciPy, ) as a function of the number of FM steps. The use of the logit-normal distribution does not appear to offer any particular advantage over the uniform distribution, in terms of either residual error or distribution. In contrast, switching to a stochastic sampler reduces the residual error if the number of function evaluations exceeds 10. However, this improvement diminishes when the number of FM steps becomes too large, as the distributional quality worsens. Particularly, with 20 integration steps, the proposed sampler reduces the residual error by a significant factor of 3×3$\times$3 ×, while the Wasserstein distance is only 30% lower for the deterministic sampler. Further results showing pressure and permeability samples and an additional UNet are reported in Figure 10 and 11 of the appendix.
@@ 4.2 Kolmogorov flow
Kolmogorov flow differs from the Darcy case as it is conditioned on the Reynolds number, and in this case the physical residual only matches parts of the underlying PDE. Here, the PBFM model reduces the average residual MSE by a factor of 1.68×1.68$\times$1.68 × compared to the baseline, while the PBFM-1 approach without unrolling offers only limited improvement of 1.05×1.05$\times$1.05 ×, as shown Figure 6. This difference is clearly reflected in the panel on the left, which visually depict the residual MSE. When examining the velocity distributions, all FM variants yield nearly overlapping profiles, although none fully capture the extrema of the reference data. In the appendix, Figure 13 and 14 report the instantaneous velocity fields along the divergence and an example of mean and standard deviation, respectively. This test case highlights the importance of accurate samples via unrolling to fully leverage the physics residuals at training time.
@@ 4.3 Dynamic Stall
Figure 7 summarizes the key results, including the residual MSE, Wasserstein distance. Focusing first on the residual error, we observe a consistent improvement when progressing from the FM baseline to PBFM, with the PBFM-1 providing a 2×2$\times$2 × reduction compared to the baseline. The incorporation of residual loss and unrolling not only reduces the average error of a 8×8$\times$8 × factor but also sharpens the residual error distribution, concentrating values closer to the mean. In terms of data distribution, the Wasserstein distance reveals that while the PBFM-1 framework alone may degrade distributional accuracy, the addition of unrolling effectively regularizes the predictions, narrowing the gap with the ground truth distribution.
@@ 6 Conclusions
In this paper, we introduce PBFM, a novel generative model designed to improve physical consistency while preserving the strengths of flow matching approach for high-dimensional data generation. Our framework provides a principled mechanism to minimize physical residuals, arising from PDEs or algebraic constraints, in a conflict-free manner. We conduct extensive benchmarks across three representative physical systems, demonstrating the versatility and robustness of our approach. Notably, PBFM achieves up to an 8$\times$ reduction in physical residuals while accurately reproducing target data distributions. The incorporation of temporal unrolling plays a key role, enabling improved final state approximations and mitigating localized residual peaks in challenging regions of the solution space.
@@ B.2 Kolmogorov flow
We generate two distinct datasets (training and validation) for the Kolmogorov flow problem with Reynolds numbers in the range [100, 500]100500[100,\ 500][ 100 , 500 ], using a spatial resolution of $128\times 128$. The simulation is based on the vorticity–stream function formulation. The velocity field is obtained from the computed vorticity through the stream function, while the pressure field is derived by solving the pressure Poisson equation. We employ TorchFSM (Liu et al., 2025b) to perform GPU-accelerated flow simulations using the Fourier spectral method. The training dataset includes 32 different flow conditions sampled via a Halton sequence (Kocis and Whiten, 1997), while the validation dataset contains 16 conditions. For each condition, 256 simulations are conducted with slightly perturbed initial states. Simulations are run for 10 000 time steps to reach a statistically steady state, followed by data sampling every 4 000 time steps. With a time step size of $dt=1/\text{Re}$, this yields 1 024 snapshots per condition.
@@ C.3 Dynamic stall
Figure 17 compares the performance of the proposed models as a function of the number of FM integration steps. The first panel focuses on the physical residual MSE, indicating that the best performance is achieved with 10 FM steps, with PBFM consistently delivering an 8×8$\times$8 × reduction compared to the baseline. The second panel reports the Wasserstein distance, showing that PBFM also narrows the gap with the reference data distribution, reaching its optimal value at 20 FM steps. Finally, the MSE of the mean and standard deviation reveals that the proposed methods preserve the already strong performance of the baseline, introducing only minor deviations.
@@ Appendix F Architecture and training details
Learning rate $3\cdot 10^{-5}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$
@@ Appendix G Performance evaluation
1 $3.23\cdot 10^{-3}$ $3.38\cdot 10^{-3}$ $3.80\cdot 10^{-3}$ $2.27\cdot 10^{0}$ $6.16\cdot 10^{0}$ $1.20\cdot 10^{1}$
@@
50 $3.21\cdot 10^{-1}$ $5.70\cdot 10^{-1}$ $1.06\cdot 10^{0}$ $1.12\cdot 10^{2}$ $3.08\cdot 10^{2}$ $5.99\cdot 10^{2}$
@@ Appendix H Sampler implementation
$dt\leftarrow 1/n$triangleright $n$ is the number of integration steps
@@
         $x_{t}\leftarrow x_{t}+(1-t)\cdot u_{t}^{\theta}$triangleright Integrate to $t=1$
@@
         $x_{t}\leftarrow(1-t)\cdot\mathcal{N}(0,1)+t\cdot x_{t}$triangleright Return to $t+dt$ using new generated normal noise
@@
         $x_{t}\leftarrow x_{t}+dt\cdot u_{t}^{\theta}$triangleright Standard deterministic sampler, stochasticity is embedded in $x_{0}$

### v2

@@ Abstract
Generative machine learning methods, such as diffusion models and flow matching, have shown great potential in modeling complex system behaviors and building efficient surrogate models. However, these methods typically learn the underlying physics implicitly from data. We propose Physics-Based Flow Matching (PBFM), a novel generative framework that explicitly embeds physical constraints, both PDE residuals and algebraic relations, into the flow matching objective. We also introduce temporal unrolling at training time that improves the accuracy of the final, noise-free sample prediction. Our method jointly minimizes the flow matching loss and the physics-based residual loss without requiring hyperparameter tuning of their relative weights. Additionally, we analyze the role of the minimum noise level, $\sigma_{\min}$, in the context of physical constraints and evaluate a stochastic sampling strategy that helps to reduce physical residuals. Through extensive benchmarks on three representative PDE problems, we show that our approach yields up to an $8\times$ more accurate physical residuals compared to FM, while clearly outperforming existing algorithms in terms of distributional accuracy. PBFM thus provides a principled and efficient framework for surrogate modeling, uncertainty quantification, and accelerated simulation in physics and engineering applications. Source code and datasets are available at https://github.com/tum-pbs/PBFM.
@@ 2.3 Improving physical and distributional accuracy
An important requirement for generative models in the context of physics applications is the necessity to produce high-accuracy samples that are on-par with traditional methods. We identify the reconstruction of the final, noise-free, sample during training as a crucial aspect for accurately evaluating the physical residual. The integration of the intermediate sample at time $t$ forward to the final time is performed with ODE integration (Lipman et al., 2024). Employing the predicted velocity field $u_{t}^{\theta}$ by integrating over a single time step of size 1t1-t. While trajectories are straight in theory, relying on just one integration step limits prediction accuracy, particularly when $t$ is close to zero. To improve the final prediction, we incorporate unrolling (Monga et al., 2021), by performing $n$ ODE integration steps of size (1t)/n(1-t)/n. The integration over multiple time steps has an averaging effect, which naturally promotes trajectories that are closer to the theoretically straight paths. Unrolling is applied via a curriculum, increasing the number of unrolling steps over the course of training. We additionally down-weigh less accurate predictions near $t=0$ by applying a scaling factor of $t^{p},\ p_{\text{opt}}=1$ in the residual loss computation. A detailed analysis of different power laws is provided in Appendix D. The more accurate prediction of $x_{1}$ substantially improves the evaluation of the physics residual in eq. 2, which in turn yields a better learning direction.
@@
$dt\leftarrow(1-t)/n$ $\triangleright$ Compute dtdt to reach the final state at $t=1$
$\widetilde{t}\leftarrow t$ $\triangleright$ Keep the starting time $t$ to weight the residual loss
$u_{t}^{\theta}\leftarrow\text{model}(x_{t},\ t)$ $\triangleright$ Flow matching velocity $u_{t}^{\theta}$ is not part of unrolling
@@
[…]
  $\widetilde{x}_{1}\leftarrow\widetilde{x}_{1}+dt\cdot\widetilde{u}_{t}^{\theta}$ $\triangleright$ Current fields $\widetilde{x}_{1}$ are updated until $t=1$
@@
[…]
$\mathcal{L}_{\mathcal{R}}\leftarrow\|t^{p}\cdot\mathcal{R}\|_{2}$ $\triangleright$ Residuals are weighted with $t^{p},\ p_{\text{opt}}=1$
@@
$\nabla_{\theta}$ $\leftarrow$ compute $\mathbf{g}_{\text{update}}$ via Eq. 3 $\triangleright$ Conflict-free update
@@ 3 Experimental Setup
The second benchmark is the two-dimensional Kolmogorov flow over a $128\times 128$ spatial domain with periodic boundary conditions. The dataset is generated using a spectral solver and consists of velocity field snapshots for various Reynolds numbers in the range [100; 500][100;\ 500]. The Reynolds number conditions the data generation, influencing the turbulence scales and flow complexity. The training set includes data for 32 Reynolds numbers with 1024 temporal snapshots each. The test set additionally includes 16 unseen Reynolds numbers. The data distribution reflects the temporal variation within each flow regime. The prediction target consists of the two velocity components, which are expected to satisfy the conservation of mass via:
@@ 4.1 Darcy flow
We begin by evaluating overall performance, focusing on residual mean absolute error (MAE), along with pressure and permeability distributions. The results are summarized in Figure 3. The first panel illustrates how residual error evolves with the number of unrolling steps used during training. The bar plot shows that MAE decreases significantly from approximately 1.7 for FM to about 0.6 for PBFM. Notably, the improvement achieved by introducing the additional loss term is comparable in magnitude to the improvement gained via unrolling. When comparing across methods, we observe that the DM baseline performs similarly to FM. CoCoGen produces an $\approx\!10\%$ lower residual error than FM but ca. $20\%$ higher than PBFM-1, at the same time incurring a higher compute cost for inference. While PIDM-ME yields the lowest residual error, this performance comes at a steep cost: As shown in the last two panels, the residual scores for PIDM-ME are associated with a severely degraded pressure distribution, characterized by values constrained within the narrow range $\pm 0.2$. This artificially reduces the residual due to the lower magnitude of $P$, rather than improved fidelity. In comparison, all variants of our framework more faithfully reproduce the true pressure distribution. Regarding permeability, the DM baseline, as well as CoCoGen, exhibit a loss of distributional fidelity, particularly at the extrema. While our framework also shows a slight degradation for the least probable values, it successfully recovers the maximum $K$ values around 30, which are approximately six orders of magnitude less frequent than the central mode of the distribution. Figure 4 visualizes the trend of decreasing residual error as the FM model is progressively improved.
@@
To complete the analysis on the Darcy dataset, we evaluated the PBFM-1 framework for sampling time $t$ during training with the logit-normal distribution, comparing it to the standard uniform distribution. We also analyze the effect of different samplers: deterministic and stochastic. Figure 5 reports the residual error along with the Wasserstein distance (Panaretos and Zemel, 2019; SciPy, ) as a function of the number of FM steps. The use of the logit-normal distribution does not appear to offer any particular advantage over the uniform distribution, in terms of either residual error or distribution. In contrast, switching to a stochastic sampler reduces the residual error if the number of function evaluations exceeds 10. However, this improvement diminishes when the number of FM steps becomes too large, as the distributional quality worsens. Particularly, with 20 integration steps, the proposed sampler reduces the residual error by a significant factor of $3\times$, while the Wasserstein distance is only 30% lower for the deterministic sampler. Further results showing pressure and permeability samples and an additional UNet are reported in Figure 10 and 11 of the appendix.
@@ 4.2 Kolmogorov flow
Kolmogorov flow differs from the Darcy case as it is conditioned on the Reynolds number, and in this case the physical residual only matches parts of the underlying PDE. Here, the PBFM model reduces the average residual MSE by a factor of $1.68\times$ compared to the baseline, while the PBFM-1 approach without unrolling offers only limited improvement of $1.05\times$, as shown Figure 6. This difference is clearly reflected in the panel on the left, which visually depict the residual MSE. When examining the velocity distributions, all FM variants yield nearly overlapping profiles, although none fully capture the extrema of the reference data. In the appendix, Figure 13 and 14 report the instantaneous velocity fields along the divergence and an example of mean and standard deviation, respectively. This test case highlights the importance of accurate samples via unrolling to fully leverage the physics residuals at training time.
@@ 4.3 Dynamic Stall
Figure 7 summarizes the key results, including the residual MSE, Wasserstein distance. Focusing first on the residual error, we observe a consistent improvement when progressing from the FM baseline to PBFM, with the PBFM-1 providing a $2\times$ reduction compared to the baseline. The incorporation of residual loss and unrolling not only reduces the average error of a $8\times$ factor but also sharpens the residual error distribution, concentrating values closer to the mean. In terms of data distribution, the Wasserstein distance reveals that while the PBFM-1 framework alone may degrade distributional accuracy, the addition of unrolling effectively regularizes the predictions, narrowing the gap with the ground truth distribution.
@@ 6 Conclusions
[…]
@@ B.2 Kolmogorov flow
We generate two distinct datasets (training and validation) for the Kolmogorov flow problem with Reynolds numbers in the range [100, 500][100,\ 500], using a spatial resolution of $128\times 128$. The simulation is based on the vorticity–stream function formulation. The velocity field is obtained from the computed vorticity through the stream function, while the pressure field is derived by solving the pressure Poisson equation. We employ TorchFSM (Liu et al., 2025b) to perform GPU-accelerated flow simulations using the Fourier spectral method. The training dataset includes 32 different flow conditions sampled via a Halton sequence (Kocis and Whiten, 1997), while the validation dataset contains 16 conditions. For each condition, 256 simulations are conducted with slightly perturbed initial states. Simulations are run for 10 000 time steps to reach a statistically steady state, followed by data sampling every 4 000 time steps. With a time step size of $dt=1/\text{Re}$, this yields 1 024 snapshots per condition.
@@ C.3 Dynamic stall
Figure 17 compares the performance of the proposed models as a function of the number of FM integration steps. The first panel focuses on the physical residual MSE, indicating that the best performance is achieved with 10 FM steps, with PBFM consistently delivering an $8\times$ reduction compared to the baseline. The second panel reports the Wasserstein distance, showing that PBFM also narrows the gap with the reference data distribution, reaching its optimal value at 20 FM steps. Finally, the MSE of the mean and standard deviation reveals that the proposed methods preserve the already strong performance of the baseline, introducing only minor deviations.
@@ Appendix F Architecture and training details
Learning rate $3\cdot 10^{-5}$ $\cdot 10^{-4} 11$$\cdot 10^{-4}$
@@ Appendix G Performance evaluation
1 $23\cdot 10^{-3} 3.3$$38\cdot 10^{-3}$ $3.80\cdot 10^{-3}$ $2.27\cdot 10^{0}$ $6.16\cdot 10^{0}$ $1.20\cdot 10^{1}$
@@
50 $3.21\cdot 10^{-1}$ $5.70\cdot 10^{-1}$ $06\cdot 10^{0} 1.1$$12\cdot 10^{2}$ $3.08\cdot 10^{2}$ $5.99\cdot 10^{2}$
@@ Appendix H Sampler implementation
$dt\leftarrow 1/n$ $\triangleright$ $n$ is the number of integration steps
@@
   $x_{t}\leftarrow x_{t}+(1-t)\cdot u_{t}^{\theta}$ $\triangleright$ Integrate to $t=1$
@@
   $x_{t}\leftarrow(1-t)\cdot\mathcal{N}(0,1)+t\cdot x_{t}$ $\triangleright$ Return to $t+dt$ using new generated normal noise
@@
   $x_{t}\leftarrow x_{t}+dt\cdot u_{t}^{\theta}$ $\triangleright$ Standard deterministic sampler, stochasticity is embedded in $x_{0}$